x+2x^1/2-15=0
How is this solve by first writing it as a quadratic equation?

Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

It's a quadratic in x^1/2
Sub y for x^1/2
--> y^2 + 2y - 15 = 0

Answer by ewatrrr(24785)


Hi,
x+2x^1/2-15=0
solve by first writing it as a quadratic equation
2x^1/2= 15-x
squaring both sides of the equation
4x = (15-x)(15-x)
4x = 225 - 30x + x^2
x^2 -34x + 225 = 0
factoring
(x-25)(x-9)= 0
(x-25)=0
x = 25
(x-9)= 0
x = 9